write the linear equations of the line joining the points (0,70)and(14,70)
Found 2 solutions by jim_thompson5910, lwsshak3:
Answer by jim_thompson5910(35256)

Both points have a y coordinate of 70, so the equation is simply y+=+70. This is a line where every point has a y coordinate of 70.

Answer by lwsshak3(11628)

write the linear equations of the line joining the points (0,70)and(14,70)
equation of a line: y=mx+b, m=slope,b=y-intercept
slope=∆y/∆x=(70-70)/(14-0)=0/14=0
y-intercept=70
equation: y=0*x+70
or y=70, just a horizontal line